# SEE: See Everything Every Time - Adaptive Brightness Adjustment for Broad Light Range Images via Events

### v1

@@ Abstract
Figure 1: (a) and (b): Brightness distributions of the SDE dataset (0$\sim$0.45, low to normal light) and our SEE-600K dataset (0$\sim$1, a broader light range). (c): Previous methods [1, 2] directly map low-light images to normal-light images. (d): Our SEENet accepts inputs across a broader brightness range and adjusts output brightness through prompts. $f_{b}$ refers to the function that calculates the brightness of an image.
@@ 1 Introduction
Every day, from daylight to nighttime, the illuminance varies from about 100,000 lux (bright sunlight) to approximately 0.1 lux (starlight)  [3]. Maintaining stable imaging under diverse natural lighting conditions is a significant challenge that energizes many embedded applications [4]. To achieve this, a series of influential works have emerged, including automatic exposure [5], exposures correction [6], low-light enhancement [7] and high dynamic range (HDR) imaging [8]. However, traditional cameras are limited by their imaging principle of synchronously capturing intensity values across the entire sensor, with a dynamic range of only 60 to 80 $dB$  [9, 10]. Consequently, these traditional methods find it difficult to capture imaging information under a wide range of lighting conditions at the input [11, 12]. If the exposure is inaccurate - over and under exposures - traditional embedded cameras lose the potential to restore images under complex lighting conditions due to limited bits-width and noise. Unlike traditional cameras, event cameras  [12] asynchronously record pixel-level changes in illumination, outputting the direction of intensity change (positive or negative) at each pixel with extremely high dynamic range (120 $dB$), which far exceeds the capability of traditional cameras in capturing various lighting intensity. Research leveraging the events for image brightness enhancement can be divided into three categories. (1) event-based image reconstruction, which aims to reconstruct images only from events. However, these methods [10, 13, 14] rely solely on events, facing uncertainties during reconstruction, and the events usually contain noise, which leads to color distortion and limited capabilities of generalization. (2) event-guided HDR imaging  [15, 16, 17] , which targets to employ events to extend the dynamic range of images or videos to match human vision.  Cui et al. [15] introduced the first real-world dataset containing paired color events, low dynamic range images, and HDR images, with only includes 1,000 HDR image pairs.  Messikommer et al. [17] used nine images with different exposures to synthesize an HDR image as the ground truth and utilized multi-exposure frames and events as inputs to generate an HDR image. While HDR imaging aims to expand dynamic range, collecting HDR datasets is difficult, and these methods have not been evaluated for tasks like low-light enhancement or high-light restoration [18, 19, 1, 2, 20]. (3) event-guided low-light enhancement  [1, 21, 22, 2], which is designed to adjust low-light images to normal-light conditions through brightness adjustment and noise reduction. Liang et al. [2] represents the latest research and proposed the first event-based low-light image enhancement dataset, SDE (see Fig. 1 (a)). Prior to this, Liang et al. [1], Liu et al. [21], Jiang et al. [22] explored using motion information from events and employed varying neural networks to improve the mapping from low-light images to normal-light ones, as shown in Fig. 1 (c). However, these strategies only focus on the improvement of mapping ability for low-light inputs, limiting their capacity to adjust brightness across a broader range of lighting conditions, e.g., normal or high-light images. The fundamental reason lies in the fact that these methods rely on datasets (such as SDE) that control brightness using a neutral density filter, i.e., they create paired low-light and normal-light images with only one fixed lighting ratio between the two. This approach forces the model to learn a rigid multiplicative enhancement strategy that only works within the narrow bounds of low-light to normal-light transitions. Consequently, the methods trained on such datasets are unable to generalize well to broad lighting conditions, particularly those involving high-light scenarios or complex lighting variations. This limitation arises because the training data does not reflect the full diversity of lighting conditions encountered in real-world applications, where illumination can change dynamically and across a much wider range. Therefore, these methods lack the flexibility needed to adjust brightness across a broader range of lighting variations.
@@
Our method has been evaluated on two real-world datasets, SDE [2] and SEE-600K. Quantitative results demonstrate that our framework fits well to a broader range of lighting conditions (Sec. 6). Furthermore, our framework allows for smooth brightness adjustment, providing precise exposure control. Therefore, this flexibility significantly improves post-processing capabilities and enables potential applications in advanced imaging and processing tasks.
@@ 2 Related Works
Frame-based: These brightness enhancement methods aim to improve image quality under challenging illumination conditions. Retinexformer [23] and other Retinex-based frameworks [24, 25, 26] decompose reflectance and illumination with complex training pipelines. Other approaches, e.g., structure-aware models [27, 28], utilize edge detection or semantic-aware guidance to achieve sharper and more realistic results. Exposure correction strategies [29, 30, 31] target both overexposed and underexposed areas, leveraging multi-scale networks or perceptual image enhancement frameworks to synthesize correctly exposed images. However, the reliance on RGB frames with limited bit depth, limits the adaptability to dynamic lighting conditions, making it difficult to handle a broader range of lighting scenarios.
Event-based: These methods focus on reconstructing images or videos exclusively from event data. For instance, Duwek et al. [32] introduced a two-phase neural network combining convolutional neural networks and spiking neural networks, while Pan et al. [33] proposed the event-based double integral model to generate videos. Stoffregen et al. [13] enhanced event-based video reconstruction by introducing the new dataset. Additionally, Liu and Dragotti [34], Wang et al. [14] developed a model-based deep network to improve reconstructed video quality. However, these event-based approaches face challenges due to event data noise, often leading to color distortion and limited generalization.
Event-guided: These works are centered on enhancing images captured in low-light conditions. E.g., Zhang et al. [35] and Liu et al. [36] recovered lost details in low-light environments by reconstructing grayscale images. Similarly, Liang et al. [1] and Liu et al. [21] improved low-light video enhancement by leveraging motion information from events to enhance multi-frame videos and integrate spatiotemporal coherence. Furthermore, Jin et al. [37] and Jiang et al. [22] utilized events to recover structural details and reconstruct clear images under near-dark situations. Most notably, Liang et al. [2] introduced the first large-scale event-guided low-light enhancement dataset, which is significant for the development of this field. While these methods use events for brightness changes and structural recovery in low-light conditions, they are limited to enhancing low-light images with single mapping and cannot handle brightness adjustments across a broader range of lighting conditions, including normal- and high-light.
@@ 3 Preliminaries and New Task Definition
In this section, we formalize the physical model underlying our approach to enhance and adjust image brightness across a broader range of lighting conditions using events. Imaging is fundamentally the process of capturing the light intensity of a scene, represented as a radiance field $L(t)$ varying over a preset slot $t$. The illumination intensities of light in daily life span a vast range, from 0.1 lux (starlight) to $61e$ 6 lux (direct sunlight). The goal of brightness adjustment is to recover or estimate $L(t)$ and tone-map it into an image that is visually suitable for human perception.
@@
Traditional cameras record light signals through exposure [38]. This voltage is influenced by the Gaussian noise $N=\mathcal{N}(\mu,\sigma^{2})$ ($\mu$ is the mean and $\sigma^{2}$ is the variance), and the photon shot noise $P=\mathcal{P}(k)$, where $k\propto L(t)$ is the number of photons, proportional to light intensity. In low-light conditions, Gaussian noise becomes dominant, while in high-light conditions, photon shot noise becomes more significant. These noise types influence the final value in the RAW image, simply represented as $I_{\text{raw}}\approx\mathcal{Q}(L(t)+P+N)$, where $\mathcal{Q}$ is the quantization function that converts the continuous voltage into discrete digital signals, typically ranging from 8 to 12 bits. The shape of the image $I_{\text{raw}}$ is $H\times W\times 1$, where $H$ and $W$ are the image resolution. The RAW image is then further processed through image signal processing (ISP) $f_{\text{isp}}$, which includes multiple steps e.g., denoising, linear and non-linear transformations, resulting in a RGB image as $I_{\text{rgb}}=f_{\text{isp}}(I_{\text{raw}})$, with the shape of $H\times W\times 3$. An accurate image exposure procedure recovers $I_{\text{rgb}}$ corresponding to $L(t)$, up to a high degree meeting the following three characteristics: (1) accurate exposure: The mean value of $I_{\text{rgb}}$ falls within the range $[0.4,0.7]$ [39]. (2) noise-free: The influence of $N$ and $P$ is suppressed to a visual-acceptable level. (3) color neutrality: The gray levels calculated from the RGB channels should be consistent [40]. However, traditional cameras sometimes fail to capture sufficient details in extreme-lighting scenes. Under such low-light conditions, images may lack visible details and be contaminated by noise, while in high-light conditions, images may suffer from oversaturation, losing texture and edge information.
Event cameras asynchronously detect illumination changes at each pixel, making them ideal for capturing scenes with extreme or rapidly changing lighting conditions [12]. The event stream’s outputs are formatted as 4 components: $(x,y)$ (pixel coordinates), $t$ (timestamp), and $p\in\{+1,-1\}$ (polarity, indicating light intensity increase or decrease). Events are triggered when the change in illumination exceeds a threshold $C$ ($\Delta L=\log(L(t))-\log(L(t-\Delta t))$ where $|\Delta L|>C$). We jointly leverage the complementary information from an image $I_{rgb}$ and its corresponding events $E$ to recover a high-quality well-illuminated image $\hat{I_{\text{rgb}}}$ that accurately represents the scene radiance $L(t)$, while also allowing for adjustable brightness. To achieve this, we introduce a brightness prompt $B$ that controls the overall brightness of the output image. This allows us to map the $L(t)$ into an image that is optimally exposed for human observation. Our task setting can thus be formulated as Eq. 1, where $f_{see}$ is our proposed model.
@@
Differences with HDR Image Reconstruction: In this section, we define the brightness adjustment task based on events. It is important to note that HDR image reconstruction also involves processing both low-light and high-light images to capture more details in both dark and bright regions. However, HDR image reconstruction aims to extend the dynamic range, which is a more ambitious and challenging task. In contrast, our goal is to adjust the dynamic range, which is easier to implement and more flexible. We highlight three key differences between these two tasks: (1) Different Objectives: HDR image reconstruction aims to expand the dynamic range of an image by combining multiple exposures, capturing details in both dark and bright regions. Mathematically, this involves recovering a radiance map $R(x)$ from events and images captured at different exposure levels, represented as $R(x)=f^{-1}(I_{\text{LDR}}(x),E)$, where $I_{\text{LDR}}(x)$ is the observed LDR image. In contrast, our brightness adjustment task focuses on modifying the exposure of a single image to recover lost details, without necessarily expanding the dynamic range, as shown in Eq. 1. (2) Distinct Challenges: HDR image reconstruction seeks to transfer the HDR characteristics of events to LDR images, requiring an HDR image as supervision. In contrast, our brightness adjustment task uses event data to adjust the exposure of a single image, without needing multiple images at different exposures for supervision. (3) Different Dataset Construction: HDR reconstruction datasets require capturing multiple images at different exposure levels [17], which can be computationally expensive and difficult to handle in dynamic scenes. In contrast, our approach uses event cameras to capture pairs of images and events under varying lighting conditions, simplifying data construction and making it more adaptable to dynamic scenes.
@@
Dataset Public Available Dynamic Motion Size (Frames) Device Frames
@@
DVS-Dark [35] 17,765 DAVIS240C
LIE [22] 2,231 DAVIS346
EvLowLight [1] DAVIS346
RLED [36] 64,200 Prophesee EVK4
SDE [2] 31,477 DAVIS346
@@ 4.1 Collection Method
To achieve this collection, we propose the following solution, which consists of three key components. (1) multiple lighting conditions: Our approach is based on the principle that lighting transitions continuously from low to high intensity. Unlike previous datasets [2, 41], which captured only a single pair of low-light and normal-light conditions, we focus on multiple samples, as shown in Fig.  4. To cover a broader lighting range, we record an average of four videos per scene, using neutral density (ND) filters at three levels $(1/8,1/64,1/1000)$ and one without a filter. We also adjust the aperture and exposure settings to capture each scene under diverse lighting conditions. (2) complex motion trajectories: We employ the Universal Robots UR5e robotic arm, which can provide high stability and repeat the same non-linear trajectory with an error margin of 0.03 mm [2, 42], allowing us to capture multiple videos with spatial consistency, as exhibited in Fig. 3 (a). (3) spatio-temporal alignment: While the robotic arm guaranteed spatial alignment, asynchronous control over the camera’s start and stop times inevitably introduced timing deviations. To resolve this, we propose an IMU-based temporal alignment algorithm, as shown in Fig. 3 (b). IMU streams synchronized to events and videos with microsecond timestamps in the DVS346 camera. Additionally, the IMU stream depends only on motion trajectory and enjoys a temporal resolution of $1000$ Hz. Based on this, our algorithm achieves precise temporal alignment, ensuring synchronization across the entire dataset, as displayed in Fig. 3 (c). We introduce the IMU-based alignment method, and the alignment evaluation results as follows. Temporal IMU Registration Algorithm: We propose an IMU data registration algorithm that aligns the source sequence $S$ and target sequence $T$ by finding the optimal bias $b$ and matching length $l$ to minimize the $L_{1}$ distance between them. Given the high resolution of IMU data at $1000Hz$, an exhaustive search for the optimal bias is computationally infeasible. To address this, we introduce a multi-level iterative strategy. First, we denoise the IMU data using a Kalman filter [43]. Then, the average pooling is utilized to reduce the sequences to two additional levels, Level-1 ($S_{1}$, $T_{1}$) and Level-2 ($S_{2}$, $T_{2}$), as shown in Fig. 3 (a)-①②. This reduces computational complexity while preserving essential alignment features. The window size is chosen based on our video durations, which range from 10 to 120 seconds. We perform a coarse search for the optimal bias $b$ and matching length $l$ at the lowest resolution (Level-2). The results from this level serve as center points for finer searches at higher resolutions. Specifically, the bias and length identified at each level guide local searches at the next level up, as displayed in Fig. 3 (a)-③④⑤. At Level-1 and the original data level (Level-0), we only need to search locally around these center points. This hierarchical approach efficiently achieves high matching accuracy with significantly reduced computational effort.
Spatial-Temporal Alignment Evaluation: To evaluate the accuracy of our IMU registration algorithm, we capture the same scene twice under identical lighting conditions, as illustrated in Fig. 3 (b). We assess the alignment metric between the two image sequences by calculating the pixel-level distance at the corresponding timestamp. Alignment Metric: For each image pair, we extract keypoints using SIFT [44] and then employ the FLANN matcher [45] to find matching keypoints between the two images. Based on these matched keypoints, we compute the affine transformation matrix using RANSAC [46]. This transformation is subsequently applied to each pixel, allowing us to calculate the displacement distance for every pixel. Finally, the average pixel distance is employed as the metric for alignment. Alignment Results: In the alignment evaluation, we select scenes with well-defined textures, as illustrated in Fig. 3 (b). After calculating the pixel distances, we observe that the average pixel error between the paired images is $0.2967$ pixels. Throughout the entire time sequence, the pixel-level distance remains below $0.8$ pixels, with the majority of errors being under $0.5$ pixels, as exhibited in Fig. 3 (c). These results demonstrate that the registration accuracy of our dataset reaches sub-pixel precision.

### v2

@@ Abstract
journal: International Journal of Computer Vision Under Reviewing
[…]
@@ 1 Introduction
Every day, from daylight to nighttime, the illuminance varies from about 100,000 lux (bright sunlight) to approximately 0.1 lux (starlight)  (Koshel, 2012). Maintaining stable imaging under diverse natural lighting conditions is a significant challenge that energizes many embedded applications (Lin et al, 2024). To achieve this, a series of influential works have emerged, including automatic exposure (Bernacki, 2020), exposures correction (Yuan and Sun, 2012), low-light enhancement (Li et al, 2021) and high dynamic range (HDR) imaging (McCann and Rizzi, 2011). However, traditional cameras are limited by their imaging principle of synchronously capturing intensity values across the entire sensor (Hasinoff et al, 2016; Rebecq et al, 2019). Consequently, these traditional methods find it difficult to capture imaging information under a wide range of lighting conditions at the input (Gehrig and Scaramuzza, 2024; Gallego et al, 2020). If the exposure is inaccurate - over and under exposures - traditional embedded cameras lose the potential to restore images under complex lighting conditions due to limited bits-width and noise. Unlike traditional cameras, event cameras  (Gallego et al, 2020) asynchronously record pixel-level changes in illumination, outputting the direction of intensity change (positive or negative) at each pixel with extremely high dynamic range (120 $dB$), which far exceeds the capability of traditional cameras in capturing various lighting intensity.
@@
Research leveraging the events for image brightness enhancement can be divided into three categories. (1) event-based image reconstruction, which aims to reconstruct images only from events. However, these methods (Rebecq et al, 2019; Stoffregen et al, 2020; Wang et al, 2024) rely solely on events, facing uncertainties during reconstruction, and the events usually contain noise, which leads to color distortion and limited capabilities of generalization. (2) event-guided HDR imaging  (Cui et al, 2024; Yang et al, 2023; Messikommer et al, 2022) , which targets to employ events to extend the dynamic range of images or videos to match human vision.  Cui et al (2024) introduced the first real-world dataset containing paired color events, low dynamic range images, and HDR images, with only includes 1,000 HDR image pairs.  Messikommer et al (2022) used nine images with different exposures to synthesize an HDR image as the ground truth and utilized multi-exposure frames and events as inputs to generate an HDR image. While HDR imaging aims to expand dynamic range, collecting HDR datasets is difficult, and these methods have not been evaluated for tasks like low-light enhancement or high-light restoration (Tursun et al, 2015; Chen et al, 2024; Liang et al, 2023, 2024; Jayasuriya et al, 2023). (3) event-guided low-light enhancement  (Liang et al, 2023; Liu et al, 2023; Jiang et al, 2023; Liang et al, 2024), which is designed to adjust low-light images to normal-light conditions through brightness adjustment and noise reduction. Liang et al (2024) represents the latest research and proposed the first event-based low-light image enhancement dataset, SDE (see Fig. 1 (a)). Prior to this, Liang et al (2023); Liu et al (2023); Jiang et al (2023) explored using motion information from events and employed varying neural networks to improve the mapping from low-light images to normal-light ones, as shown in Fig. 1 (c). However, these strategies only focus on the improvement of mapping ability for low-light inputs, limiting their capacity to adjust brightness across a broader range of lighting conditions, e.g., normal or high-light images. The fundamental reason lies in the fact that these methods rely on datasets (such as SDE) that control brightness using a neutral density filter, i.e., they create paired low-light and normal-light images with only one fixed lighting ratio between the two. This approach forces the model to learn a rigid multiplicative enhancement strategy that only works within the narrow bounds of low-light to normal-light transitions. Consequently, the methods trained on such datasets are unable to generalize well to broad lighting conditions, particularly those involving high-light scenarios or complex lighting variations. This limitation arises because the training data does not reflect the full diversity of lighting conditions encountered in real-world applications, where illumination can change dynamically and across a much wider range. Therefore, these methods lack the flexibility needed to adjust brightness across a broader range of lighting variations.
@@
Our method has been evaluated on two real-world datasets, SDE (Liang et al, 2024) and SEE-600K. Quantitative results demonstrate that our framework fits well to a broader range of lighting conditions (Sec. 6). Furthermore, our framework allows for smooth brightness adjustment, providing precise exposure control. Therefore, this flexibility significantly improves post-processing capabilities and enables potential applications in advanced imaging and processing tasks.
@@ 2 Related Works
Frame-based: These brightness enhancement methods aim to improve image quality under challenging illumination conditions. Retinexformer (Cai et al, 2023) and other Retinex-based frameworks (Zhang et al, 2021; Wu et al, 2022; Fu et al, 2023) decompose reflectance and illumination with complex training pipelines. Other approaches, e.g., structure-aware models (Xu et al, 2023b; Wang et al, 2023b), utilize edge detection or semantic-aware guidance to achieve sharper and more realistic results. Exposure correction strategies (Afifi et al, 2021; Panetta et al, 2022; Ma et al, 2020) target both overexposed and underexposed areas, leveraging multi-scale networks or perceptual image enhancement frameworks to synthesize correctly exposed images. However, the reliance on RGB frames with limited bit depth, limits the adaptability to dynamic lighting conditions, making it difficult to handle a broader range of lighting scenarios.
Event-based: These methods focus on reconstructing images or videos exclusively from event data. For instance, Duwek et al (2021) introduced a two-phase neural network combining convolutional neural networks and spiking neural networks, while Pan et al (2019) proposed the event-based double integral model to generate videos. Stoffregen et al (2020) enhanced event-based video reconstruction by introducing the new dataset. Additionally, Liu and Dragotti (2023); Wang et al (2024) developed a model-based deep network to improve reconstructed video quality. However, these event-based approaches face challenges due to event data noise, often leading to color distortion and limited generalization.
Event-guided: These works are centered on enhancing images captured in low-light conditions. E.g., Zhang et al (2020) and Liu et al (2024) recovered lost details in low-light environments by reconstructing grayscale images. Similarly, Liang et al (2023) and Liu et al (2023) improved low-light video enhancement by leveraging motion information from events to enhance multi-frame videos and integrate spatiotemporal coherence. Furthermore, Jin et al (2023) and Jiang et al (2023) utilized events to recover structural details and reconstruct clear images under near-dark situations. Most notably, Liang et al (2024) introduced the first large-scale event-guided low-light enhancement dataset, which is significant for the development of this field. While these methods use events for brightness changes and structural recovery in low-light conditions, they are limited to enhancing low-light images with single mapping and cannot handle brightness adjustments across a broader range of lighting conditions, including normal- and high-light.
@@ 3 Preliminaries and New Task Definition
In this section, we formalize the physical model underlying our approach to enhance and adjust image brightness across a broader range of lighting conditions using events. Imaging is fundamentally the process of capturing the light intensity of a scene, represented as a radiance field $L(t)$ varying over a preset slot $t$. The illumination intensities of light in daily life span a vast range, from 0.1 lux (starlight) to $1e6$ lux (direct sunlight). The goal of brightness adjustment is to recover or estimate $L(t)$ and tone-map it into an image that is visually suitable for human perception.
@@
Traditional cameras record light signals through exposure (Mendis et al, 1997). This voltage is influenced by the Gaussian noise $N=\mathcal{N}(\mu,\sigma^{2})$ ($\mu$ is the mean and $\sigma^{2}$ is the variance), and the photon shot noise $P=\mathcal{P}(k)$, where $k\propto L(t)$ is the number of photons, proportional to light intensity. In low-light conditions, Gaussian noise becomes dominant, while in high-light conditions, photon shot noise becomes more significant. These noise types influence the final value in the RAW image, simply represented as $I_{\text{raw}}\approx\mathcal{Q}(L(t)+P+N)$, where $\mathcal{Q}$ is the quantization function that converts the continuous voltage into discrete digital signals, typically ranging from 8 to 12 bits. The shape of the image $I_{\text{raw}}$ is $H\times W\times 1$, where $H$ and $W$ are the image resolution. The RAW image is then further processed through image signal processing (ISP) $f_{\text{isp}}$, which includes multiple steps e.g., denoising, linear and non-linear transformations, resulting in a RGB image as $I_{\text{rgb}}=f_{\text{isp}}(I_{\text{raw}})$, with the shape of $H\times W\times 3$. An accurate image exposure procedure recovers $I_{\text{rgb}}$ corresponding to $L(t)$, up to a high degree meeting the following three characteristics: (1) accurate exposure: The mean value of $I_{\text{rgb}}$ falls within the range $[0.4,0.7]$ (Mertens et al, 2009). (2) noise-free: The influence of $N$ and $P$ is suppressed to a visual-acceptable level. (3) color neutrality: The gray levels calculated from the RGB channels should be consistent (Buchsbaum, 1980). However, traditional cameras sometimes fail to capture sufficient details in extreme-lighting scenes. Under such low-light conditions, images may lack visible details and be contaminated by noise, while in high-light conditions, images may suffer from oversaturation, losing texture and edge information.
Event cameras asynchronously detect illumination changes at each pixel, making them ideal for capturing scenes with extreme or rapidly changing lighting conditions (Gallego et al, 2020). The event stream’s outputs are formatted as 4 components: $(x,y)$ (pixel coordinates), $t$ (timestamp), and $p\in\{+1,-1\}$ (polarity, indicating light intensity increase or decrease). Events are triggered when the change in illumination exceeds a threshold $C$ ($\Delta L=\log(L(t))-\log(L(t-\Delta t))$ where $|\Delta L|>C$). We jointly leverage the complementary information from an image $I_{rgb}$ and its corresponding events $E$ to recover a high-quality well-illuminated image $\hat{I_{\text{rgb}}}$ that accurately represents the scene radiance $L(t)$, while also allowing for adjustable brightness. To achieve this, we introduce a brightness prompt $B$ that controls the overall brightness of the output image. This allows us to map the $L(t)$ into an image that is optimally exposed for human observation. Our task setting can thus be formulated as Eq. 1, where $f_{see}$ is our proposed model.
@@
Differences with HDR Image Reconstruction: In this section, we define the brightness adjustment task based on events. It is important to note that HDR image reconstruction also involves processing both low-light and high-light images to capture more details in both dark and bright regions. However, HDR image reconstruction aims to extend the dynamic range, which is a more ambitious and challenging task. In contrast, our goal is to adjust the dynamic range, which is easier to implement and more flexible. We highlight three key differences between these two tasks: (1) Different Objectives: HDR image reconstruction aims to expand the dynamic range of an image by combining multiple exposures, capturing details in both dark and bright regions. Mathematically, this involves recovering a radiance map $R(x)$ from events and images captured at different exposure levels, represented as $R(x)=f^{-1}(I_{\text{LDR}}(x),E)$, where $I_{\text{LDR}}(x)$ is the observed LDR image. In contrast, our brightness adjustment task focuses on modifying the exposure of a single image to recover lost details, without necessarily expanding the dynamic range, as shown in Eq. 1. (2) Distinct Challenges: HDR image reconstruction seeks to transfer the HDR characteristics of events to LDR images, requiring an HDR image as supervision. In contrast, our brightness adjustment task uses event data to adjust the exposure of a single image, without needing multiple images at different exposures for supervision. (3) Different Dataset Construction: HDR reconstruction datasets require capturing multiple images at different exposure levels (Messikommer et al, 2022), which can be computationally expensive and difficult to handle in dynamic scenes. In contrast, our approach uses event cameras to capture pairs of images and events under varying lighting conditions, simplifying data construction and making it more adaptable to dynamic scenes.
@@
Dataset
Public
Available
Dynamic
Motion
Size
Device Frames
@@
DVS-Dark (Zhang et al, 2020) 17,765 DAVIS240C
LIE (Jiang et al, 2023) 2,231 DAVIS346
EvLowLight (Liang et al, 2023) DAVIS346
RLED (Liu et al, 2024) 64,200 Prophesee EVK4
SDE (Liang et al, 2024) 31,477 DAVIS346
@@ 4.1 Collection Method
To achieve this collection, we propose the following solution, which consists of three key components. (1) multiple lighting conditions: Our approach is based on the principle that lighting transitions continuously from low to high intensity. Unlike previous datasets (Liang et al, 2024; Wang et al, 2021), which captured only a single pair of low-light and normal-light conditions, we focus on multiple samples, as shown in Fig.  4. To cover a broader lighting range, we record an average of four videos per scene, using neutral density (ND) filters at three levels $(1/8,1/64,1/1000)$ and one without a filter. We also adjust the aperture and exposure settings to capture each scene under diverse lighting conditions. (2) complex motion trajectories: We employ the Universal Robots UR5e robotic arm, which can provide high stability and repeat the same non-linear trajectory with an error margin of 0.03 mm (Liang et al, 2024; Brey et al, 2024), allowing us to capture multiple videos with spatial consistency, as exhibited in Fig. 3 (a). (3) spatio-temporal alignment: While the robotic arm guaranteed spatial alignment, asynchronous control over the camera’s start and stop times inevitably introduced timing deviations. To resolve this, we propose an IMU-based temporal alignment algorithm, as shown in Fig. 3 (b). IMU streams synchronized to events and videos with microsecond timestamps in the DVS346 camera. Additionally, the IMU stream depends only on motion trajectory and enjoys a temporal resolution of $1000$ Hz. Based on this, our algorithm achieves precise temporal alignment, ensuring synchronization across the entire dataset, as displayed in Fig. 3 (c). We introduce the IMU-based alignment method, and the alignment evaluation results as follows.
@@
Temporal IMU Registration Algorithm: We propose an IMU data registration algorithm that aligns the source sequence $S$ and target sequence $T$ by finding the optimal bias $b$ and matching length $l$ to minimize the $L_{1}$ distance between them. Given the high resolution of IMU data at $1000Hz$, an exhaustive search for the optimal bias is computationally infeasible. To address this, we introduce a multi-level iterative strategy. First, we denoise the IMU data using a Kalman filter (Mirzaei and Roumeliotis, 2008). Then, the average pooling is utilized to reduce the sequences to two additional levels, Level-1 ($S_{1}$, $T_{1}$) and Level-2 ($S_{2}$, $T_{2}$), as shown in Fig. 3 (a)-①②. This reduces computational complexity while preserving essential alignment features. The window size is chosen based on our video durations, which range from 10 to 120 seconds. We perform a coarse search for the optimal bias $b$ and matching length $l$ at the lowest resolution (Level-2). The results from this level serve as center points for finer searches at higher resolutions. Specifically, the bias and length identified at each level guide local searches at the next level up, as displayed in Fig. 3 (a)-③④⑤. At Level-1 and the original data level (Level-0), we only need to search locally around these center points. This hierarchical approach efficiently achieves high matching accuracy with significantly reduced computational effort.
Spatial-Temporal Alignment Evaluation: To evaluate the accuracy of our IMU registration algorithm, we capture the same scene twice under identical lighting conditions, as illustrated in Fig. 3 (b). We assess the alignment metric between the two image sequences by calculating the pixel-level distance at the corresponding timestamp. Alignment Metric: For each image pair, we extract keypoints using SIFT (Lowe, 2004) and then employ the FLANN matcher (Muja and Lowe, 2009) to find matching keypoints between the two images. Based on these matched keypoints, we compute the affine transformation matrix using RANSAC (Fischler and Bolles, 1981). This transformation is subsequently applied to each pixel, allowing us to calculate the displacement distance for every pixel. Finally, the average pixel distance is employed as the metric for alignment. Alignment Results: In the alignment evaluation, we select scenes with well-defined textures, as illustrated in Fig. 3 (b). After calculating the pixel distances, we observe that the average pixel error between the paired images is $0.2967$ pixels. Throughout the entire time sequence, the pixel-level distance remains below $0.8$ pixels, with the majority of errors being under $0.5$ pixels, as exhibited in Fig. 3 (c). These results demonstrate that the registration accuracy of our dataset reaches sub-pixel precision.